What is a point of inflexion ? Does f(x) = x3

 have a point of inflexion at x = 0?


1
Expert's answer
2021-04-06T07:25:19-0400

At point of inflexion, concavity of any function (or curve) changes i.e. the curve becomes convex upwards (or convex : f”(x)>0) to convex downwards (or concave : f”(x)<0) & vice-versa. Simply in a word, it is the point where second derivative of a function is zero.


Now for the curve "f(x)=x^3"

"f'(x)=3x^2" And. "f''(x)=6x"


And at point x=0 , f”(0)=0


Thus "f(x)=x^3" has a point of inflexion at x=0 .